OpenStudy (zamzam-hussain):

a funnel 0.24m in diameter at the top and 24mm in diameter at the bottom is 0.54m deep.what will be the total volume of funnel ? plz tell me

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

what you are gonna do is 1st - find the average of the two radius

OpenStudy (anonymous):

No, don't use the volume of a cone

OpenStudy (anonymous):

1st- find the average of the two radii 2nd - plug in the average of the two radii to the usual cylinder formula pi x r^2 x h

OpenStudy (zamzam-hussain):

k then

Parth (parthkohli):

no

OpenStudy (anonymous):

It's a truncated cone

Parth (parthkohli):

\[\rm V_{frustum } = \frac{1}{3}\pi h(r_1^2 + r_2 ^2 + r_1 r_2)\]

OpenStudy (zamzam-hussain):

oohhh parth kohli thnkx

OpenStudy (anonymous):

volume of larger cone - volume of smaller cone :D

OpenStudy (mathmate):

I thought @kliedako was going to use the universal formula: V=h(A1+4A2+A3)/6 which works for cones, prisms, cylinders, pyramids, spheres and even truncated cones. That would have given: \(\rm V_{truncated~cone}=\dfrac{\pi(r_1^2+4(\frac{r_1+r_2}{2})^2+r_2^2)h}{6}\) which could be simplified to \(\rm V_{truncated~cone} = \frac{1}{3}\pi h(r_1^2 + r_2 ^2 + r_1 r_2)\) as @parthkoli had it.